OpenStudy (anonymous):

Write the complex number in polar form with argument θ between 0 and 2π. −9πi

OpenStudy (kropot72):

r=+sqrt(a^2+b^2)=+sqrt(0^2+(-9\pi)^2=9\pi tan theta=b/a=inf theta=\pi/2 Polar form is: 9\pi(cos\pi/2+i*sin\pi/2)

OpenStudy (kropot72):

Sorry. Have to make a correction: tan theta=b/a=(-9pi)/0=-(infinity) theta=(3*pi)/2 Polar form is: 9*pi(cos((3*pi)/2) + i*sin((3*pi)/2))